# Small solutions to inhomogeneous and homogeneous quadratic congruences modulo prime powers

### v2

@@ 1. Introduction and statement of results
with $1\leq x_{i}\ll_{k}p^{1/k}$ for $i=1,...,n$. The above condition on $n$ becomes $n>12$ if $k=2$.
@@
be an inhomogeneous quadratic form with fixed integer coefficients $\lambda_{1},...,\lambda_{n+1}$ such that $(\lambda_{1}\cdots\lambda_{n+1},p)=1$ and the homogeneous quadratic form $\lambda_{1}x_{1}^{2}+\cdots+\lambda_{n}x_{n}^{2}$ is positive definite, i.e. $\lambda_{i}>0$ for $i=1,...,n$. Set $B_{p}(Q):=\sharp\mathcal{B}(Q)/p^{n-1}$, where
@@
contains at least $(p+1)/2$ elements, and hence their intersection is non-empty, which implies that $\mathcal{B}(Q)$ is non-empty. Indeed, the set $\left\{\lambda_{i}x_{i}^{2}:x_{i}\in\mathbb{F}_{p}^{\times}\right\}$ equals either the set of quadratic residues or that of quadratic non-residues modulo $p$, which both have a cardinality of $(p-1)/2$. By the Cauchy-Davenport theorem, the sum of two such sets has at least $p-2$ elements, which is greater than or equal to $(p+1)/2$ if $p\geq 5$.
@@ 2.2. Poisson summation formula
where $\hat{\Phi}:\mathbb{R}\rightarrow\mathbb{C}$ is the Fourier transform of $\Phi$, defined by
@@ 2.4. Results on Kloosterman and Salié sums
If $p\nmid b$, then $r=ord_{p}(f)=0$ and $f^{\prime}(\alpha)\equiv 0\bmod p$ if and only if $\alpha^{2}\equiv a\overline{b}\bmod p$. Now as $p|a$, we have $a\overline{b}\equiv 0\bmod p$ and so $\alpha^{2}\equiv 0\bmod p$. But $\alpha$ varies from 1 to $p-1$, so $\alpha^{2}\equiv 0\bmod p$ cannot hold. Hence if $11\leq\alpha\leq p-$ 1, then $p^{-r}f^{\prime}(\alpha)\not\equiv 0\bmod p$ and therefore $S_{\alpha}(f;p^{n})=0$, resulting in $K_{0}(a,b,c)=0$.
@@ 2.5. Weigthed representations by quadratic forms
if $20<\epsilon\leq y/$ 2. From this, we deduce the estimate
@@
The limit on the right-hand side equals the $(n-1)$-dimensional volume of the hypersurface
@@
Since $F^{(0)}({\bf x})$ is a positive definite quadratic form, this volume is finite and proportional to $y^{(n-1)/2}$. Now taking $A:=(n-1)/2$, it follows that
@@ 3.3.1. Evaluation of Gauss sums
Suppose that $(k_{1},k_{2},...,k_{n},p^{m})=p^{r}$. We first note that using the rapid decay of $\hat{\Phi}$ and ${\bf k}\not=0$, the contributions of $r=m-1$ and $r=m$ to the right-hand side of (3.2) are negligble and hence bounded by $O(1)$ if we assume that $N\geq p^{m\varepsilon}$ and $m$ tends to infinity. So we left with the contributions of $r\in\{0,...,m-2\}$.
@@
where $G(a,b,c)$ is the quadratic Gauss sum defined in the section 2. Suppose $20\leq r\leq m-$ 2 and let $j\in\{1,...,n\}$ be such that $(k_{j},p^{m})=p^{r}$. We consider the following three cases.
@@
where $h=p^{r}h^{\prime}$ and $(h^{\prime},p)=1$. It follows that if $(k_{1},k_{2},...,k_{n},p^{m})=p^{r}$ for some $20\leq r\leq m-$ 2, then
@@
If $u=m-1,m$, similar calculations as in Case 3 give that
@@ 3.3.2. Evaluation of Kloosterman and Salié sums
with $\Delta_{j}:=\Delta/\lambda_{j}$ for $j=1,2,...,n+1$. We note that $(A,p)=1$ if and only if $(\tilde{Q}({\bf l}),p)=1$. For $k\in\mathbb{Z}$, we define
@@
Case 2(i): $n$ even, $r\leq m-2$ and $\epsilon_{p^{m-r}}=1$. In this case, we necessarily have $p\equiv 1\bmod{4}$ or $p\equiv 3\bmod{4}$ and $m-r$ even, and hence
@@
Case 2(ii): $n$ even, $r\leq m-2$ and $\epsilon_{p^{m-r}}=i$. In this case, we necessarily have $p\equiv 3\bmod{4}$ and $m-r$ odd and hence
@@ 3.3.3. Evaluation of sums involving square roots modulo prime powers
$\sum\limits_{\begin{subarray}{c}w\bmod p^{t}\\ (w,p)=1\end{subarray}}\ \sum\limits_{\begin{subarray}{c}\alpha\in\mathbb{Z}\\ \alpha\in I_{w}\end{subarray}}\left|e_{p^{t}}\left(\overline{2w}\Lambda\tilde{c}\alpha\right)\right|\ll\sum\limits_{\begin{subarray}{c}w\bmod p^{t}\\ (w,p)=1\end{subarray}}\left|\left|\frac{\overline{2w}\Lambda\tilde{c}}{p^{t}}\right|\right|^{-1}\\ =\sum\limits_{\begin{subarray}{c}\tilde{w}\bmod p^{t}\\ (\tilde{w},p)=1\end{subarray}}\left|\left|\frac{\tilde{w}}{p^{t}}\right|\right|^{-1}\ll p^{t}\log p^{t}\ll p^{s/2}\log p^{s},$
where for $x\in\mathbb{R}$, $||x||$ is the distance of $x$ to the nearest integer.
@@ 3.3.4. Final evaluation
Let $p$ be a fixed odd prime, $\mu\in\{0,1\}$, $\Lambda\in\mathbb{Z}$ with $(\Lambda,p)=1$, $r\in\mathbb{N}$ with $20\leq r\leq m-$ 2 and $1\leq K,N\leq p^{m-r}$. Define $\tau_{n}(k)$ as in (3.10) and set $s:=m-r$. Then if $n\geq 6$, we have

### v3

@@ 1. Introduction and statement of results
with $1\leq x_{i}\ll_{k}p^{1/k}$ for i=1,,ni=1,...,n. The above condition on $n$ becomes $n>12$ if $k=2$.
@@
be an inhomogeneous quadratic form with fixed integer coefficients $\lambda_{1},...,\lambda_{n+1}$ such that $(\lambda_{1}\cdots\lambda_{n+1},p)=1$ and the homogeneous quadratic form $\lambda_{1}x_{1}^{2}+\cdots+\lambda_{n}x_{n}^{2}$ is positive definite, i.e. $\lambda_{i}>0$ for i=1,,ni=1,...,n. Set $B_{p}(Q):=\sharp\mathcal{B}(Q)/p^{n-1}$, where
@@
contains at least $(p+1)/2$ elements, and hence their intersection is non-empty, which implies that $\mathcal{B}(Q)$ is non-empty. Indeed, the set $\left\{\lambda_{i}x_{i}^{2}:x_{i}\in\mathbb{F}_{p}^{\times}\right\}$ equals either the set of quadratic residues or that of quadratic non-residues modulo $p$, which both have a cardinality of (p1)/2(p-1)/2. By the Cauchy-Davenport theorem, the sum of two such sets has at least p2p-2 elements, which is greater than or equal to $(p+1)/2$ if $p\geq 5$.
@@
Data availability statement: This manuscript has no associated data.
Conflict of interest statement: The authors have no conflicts of interest to declare. All co-authors have seen and agree with the contents of the manuscript and there is no financial interest to report. We certify that the submission is original work and is not under review at any other publication.
@@ 2.2. Poisson summation formula
where $\hat{\Phi}$$\mathbb{R}\rightarrow\mathbb{C}$ is the Fourier transform of $\Phi$, defined by
@@ 2.4. Results on Kloosterman and Salié sums
If $p\nmid b$, then $r=ord_{p}(f)=0$ and $f^{\prime}(\alpha)\equiv 0\bmod p$ if and only if $\alpha^{2}\equiv a\overline{b}\bmod p$. Now as $p|a$, we have $a\overline{b}\equiv 0\bmod p$ and so $\alpha^{2}\equiv 0\bmod p$. But $\alpha$ varies from 1 to p1p-1, so $\alpha^{2}\equiv 0\bmod p$ cannot hold. Hence if $1\leq\alpha\leq p-1$, then $p^{-r}f^{\prime}(\alpha)\not\equiv 0\bmod p$ and therefore $S_{\alpha}(f;p^{n})=0$, resulting in $K_{0}(a,b,c)=0$.
@@ 2.5. Weigthed representations by quadratic forms
if $0<\epsilon\leq y/2$. From this, we deduce the estimate
@@
The limit on the right-hand side equals the (n1)(n-1)-dimensional volume of the hypersurface
@@
Since $F^{(0)}({\bf x})$ is a positive definite quadratic form, this volume is finite and proportional to $y^{(n-1)/2}$. Now taking A:=(n1)/2A:=(n-1)/2, it follows that
@@ 3.3.1. Evaluation of Gauss sums
Suppose that $(k_{1},k_{2},...,k_{n},p^{m})=p^{r}$. We first note that using the rapid decay of $\hat{\Phi}$ and ${\bf k}\not=0$, the contributions of r=m1r=m-1 and $r=m$ to the right-hand side of (3.2) are negligble and hence bounded by $O(1)$ if we assume that $N\geq p^{m\varepsilon}$ and $m$ tends to infinity. So we left with the contributions of $r\in\{0,...,m-2\}$.
@@
where $G(a,b,c)$ is the quadratic Gauss sum defined in the section 2. Suppose $0\leq r\leq m-2$ and let $j\in\{1,...,n\}$ be such that $(k_{j},p^{m})=p^{r}$. We consider the following three cases.
@@
where $h=p^{r}h^{\prime}$ and $(h^{\prime},p)=1$. It follows that if $(k_{1},k_{2},...,k_{n},p^{m})=p^{r}$ for some $0\leq r\leq m-2$, then
@@
If u=m1,mu=m-1,m, similar calculations as in Case 3 give that
@@ 3.3.2. Evaluation of Kloosterman and Salié sums
with $\Delta_{j}:=\Delta/\lambda_{j}$ for j=1,2,,n+1j=1,2,...,n+1. We note that $(A,p)=1$ if and only if $(\tilde{Q}({\bf l}),p)=1$. For $k\in\mathbb{Z}$, we define
@@
Case 2(i): $n$ even, $r\leq m-2$ and $\epsilon_{p^{m-r}}=1$. In this case, we necessarily have $p\equiv 1\bmod{4}$ or $p\equiv 3\bmod{4}$ and mrm-r even, and hence
@@
Case 2(ii): $n$ even, $r\leq m-2$ and $\epsilon_{p^{m-r}}=i$. In this case, we necessarily have $p\equiv 3\bmod{4}$ and mrm-r odd and hence
@@ 3.3.3. Evaluation of sums involving square roots modulo prime powers
[…]
where for $x\in\mathbb{R}$, $x||$ is the distance of $x$ to the nearest integer.
@@ 3.3.4. Final evaluation
Let $p$ be a fixed odd prime, $\mu\in\{0,1\}$, $\Lambda\in\mathbb{Z}$ with $(\Lambda,p)=1$, $r\in\mathbb{N}$ with $0\leq r\leq m-2$ and $1\leq K,N\leq p^{m-r}$. Define $\tau_{n}(k)$ as in (3.10) and set s:=mrs:=m-r. Then if $n\geq 6$, we have